A stack of neodymium button magnets (12 mm diameter, 75 mm tall) dropped from 0.200 m high, generates an average +2.375 Volt for 0.040 seconds, as it enters a lab coil that has 800 turns spread along 40 mm tall, each with 0.112 m circumference (28 mm long × 28 mm wide).

[a) how much B pierces the coil Area after the magnet has exited?]

d) if the stack takes 0.0375 s to exit the coil (downward), what average voltage would be generated?
Answer #1

— О: О12 м) 4 - 12 mm (= a = А - Y- п/о-оос ? = 3стхо бо О: О12 2. - о: ОО 6 т. 2 (a) . х. Дә BA at + 2. 325 - в (3. (пхо 5 )